# Offline changepoint localization using a matrix of conformal p-values

### v4

@@ Appendix E Proofs of main results
Note that $\lim_{y\uparrow F_{X}^{-1}(u)}F_{X}(y)<u\leq F_{X}(F_{X}^{-1}(u))$ by definition, so since $V\sim\mathrm{Unif}(0,1)$ we obtain
@@
Finally, for all $x>F_{X}^{-1}(u)$, we have $F_{X}(x)\geq\lim_{y\uparrow F_{X}^{-1}(u)}F_{X}(y)>u$. This means that whenever $X>F_{X}^{-1}(u)$, we have
$U=\lim_{y\uparrow X}F_{X}(y)+V\left(F_{X}(X)-\lim_{y\uparrow X}F_{X}(y)\right)>u.$
We have shown $\mathbb{P}(U\leq u,\,X>F_{X}^{-1}(u))=0$, leading to the final estimate $\mathbb{P}(U\leq u)=u$, thereby showing that $U\sim\mathrm{Unif}(0,1)$ as desired. ∎
Using Lemma E.1, we can prove coverage guarantees for our confidence sets.
@@
Integrating over $\llbracket X_{1},\dots,X_{r}\rrbracket$ shows that $p_{r}^{(t)}$ is an unconditional p-value for $\mathcal{H}_{0t}$ such that $\{p_{r}^{(t)}\}_{r=1}^{t}$ are independent [29, Proposition 3.1]. Of course, this means that under the null hypothesis, the discrepancy scores $((W_{t}^{(i)})_{i=0}^{1})_{t=1}^{n-1}$ are distributed as if the data was all i.i.d. from $\mathrm{Unif}(0,1)$. Another application of the randomized probability integral transform (Lemma E.1) gives that $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ are exactly uniformly distributed, and any of the methods in Section 4.2 will yield a valid confidence interval by the Neyman construction. ∎
Proof of Theorem C.1.
The proof of this fact is completely analogous to the proof of Theorem 3.1, except that the test used is only asymptotically valid as $t\to\infty$ and $n-t\to\infty$ simultaneously. Letting $\varphi_{t}=B_{t}-tB_{1}$ (where $(B_{t})_{t\geq 0}$ denotes a standard Brownian motion on $\mathbb{R}$), the validity of the test follows from the following theorem of [7]:
@@
$\mathbb{E}[T_{n}[s]]$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\mathbb{P}(s(X_{i})>s(X_{n+1}))+\frac{1}{2}\,\mathbb{P}(s(X_{i})=s(X_{n+1}))\right)$
The second term does not depend on the choice of $s$, so it suffices to optimize over the first term. By the randomized probability integral transform (Lemma E.1), we obtain
$\mathbb{P}(s(X_{1})>s(X_{n+1}))+\frac{1}{2}\,\mathbb{P}(s(X_{1})=s(X_{n+1}))$
@@
$F_{Y}^{*}(s(X_{1})):=\lim_{y\uparrow s(X_{1})}\,F_{Y}(y)+\theta_{1}\left(F_{Y}(s(X_{1}))-\lim_{y\uparrow s(X_{1})}\,F_{Y}(y)\right).$
@@
Fix $u\in(0,1)$ and suppose we want to test $\tilde{\mathcal{H}}_{0}:X_{1}\sim Q$ against $\tilde{\mathcal{H}}_{1}:X_{1}\sim R$, rejecting $\tilde{\mathcal{H}}_{0}$ whenever $F_{s(X_{n+1})}^{*}(s(X_{1}))\leq u$. Under $\tilde{\mathcal{H}}_{0}$, the cdf of $s(X_{1})$ is exactly $F_{s(X_{n+1})}$, so the randomized probability integral transform (Lemma E.1) gives that the Type I error of our test is $u$:
@@
The power of our test is $\mathbb{P}_{\tilde{\mathcal{H}_{1}}}\left(F_{s(X_{n+1})}^{*}(s(X_{1}))<u\right)$. By the generalized Neyman-Pearson lemma (see, for example, Theorem 3.2.1 and Section 5.9 of [13]), the power is maximized by the choice $s=s^{*}$. Integrating the bound from 0 to 1 gives the desired result. ∎
Appendix F The log-optimal e-variable uses the likelihood ratio score
While the reasoning in Theorem 5.1 yields that conformal p-values are optimized by the likelihood ratio score, it turns out to also be true that conformal e-values are also optimized by the same score. Indeed, in order to test exchangeability of $X_{1},\dots,X_{n+1}$ against the alternative that their joint distribution is given by $R^{n}\times Q$, one can show [17, Section 6.7.7] that the log-optimal e-variable is given by
@@ Appendix F The log-optimal e-variable uses the likelihood ratio score
Appendix G Convolutional neural network architecture
@@
Appendix H Additional methods for choosing the score function

### v5

@@ Appendix F Proofs of main results
Note that $\lim_{y\uparrow F_{X}^{-1}(u)}F_{X}(y)\leq u\leq F_{X}(F_{X}^{-1}(u))$ by definition, so since $V\sim\mathrm{Unif}(0,1)$ we obtain
@@
Finally, consider the case $X>F_{X}^{-1}(u)$ in Equation 1. By monotonicity of the cdf, we have
$\lim_{y\uparrow X}\,F_{X}(y)\geq F_{X}(F_{X}^{-1}(u)).$
By definition, $F_{X}(F_{X}^{-1}(u))\geq u$, so we see that
$U=\lim_{y\uparrow X}F_{X}(y)+V\left(F_{X}(X)-\lim_{y\uparrow X}F_{X}(y)\right)\geq\lim_{y\uparrow X}F_{X}(y)\geq u.$
Thus, $\{U\leq u,\,X>F_{X}^{-1}(u)\}$ can only happen if $U=u$. Except for the zero-probability event where $V=0$, this also forces $X\in S\coloneqq\{x\in\mathbb{R}:\lim_{y\uparrow x}F_{X}(y)=F_{X}(x)=u\}$ and because $F_{X}$ is constant on $S$, we find that
$\mathbb{P}(U\leq u,\,X>F_{X}^{-1}(u))\leq\mathbb{P}(X\in S)=0.$
We deduce the result from summing the three terms in Equation 1. ∎
Using Lemma F.1, we can prove coverage guarantees for our confidence sets.
@@
Integrating over $\llbracket X_{1},\dots,X_{r}\rrbracket$ shows that $p_{r}^{(t)}$ is an unconditional p-value for $\mathcal{H}_{0t}$ such that $\{p_{r}^{(t)}\}_{r=1}^{t}$ are independent [vovk2003testing, Proposition 3.1]. Of course, this means that under the null hypothesis, the discrepancy scores $((W_{t}^{(i)})_{i=0}^{1})_{t=1}^{n-1}$ are distributed as if the data was all i.i.d. from $\mathrm{Unif}(0,1)$. Another application of the randomized probability integral transform (Lemma F.1) gives that $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ are exactly uniformly distributed, and any of the methods in Section 4.2 will yield a valid confidence interval by the Neyman construction. ∎
Proof of Theorem D.1.
The proof of this fact is completely analogous to the proof of Theorem 4.1, except that the test used is only asymptotically valid as $t\to\infty$ and $n-t\to\infty$ simultaneously. Letting $\varphi_{t}=B_{t}-tB_{1}$ (where $(B_{t})_{t\geq 0}$ denotes a standard Brownian motion on $\mathbb{R}$), the validity of the test follows from the following theorem of [donsker1951invariance]:
@@
$\mathbb{E}[T_{n}[s]]$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\mathbb{P}(s(X_{i})<s(X_{n+1}))+\frac{1}{2}\,\mathbb{P}(s(X_{i})=s(X_{n+1}))\right).$
The second term does not depend on the choice of $s$, so it suffices to optimize over the first term. By the randomized probability integral transform (Lemma F.1), we obtain
$\mathbb{P}(s(X_{1})<s(X_{n+1}))+\frac{1}{2}\,\mathbb{P}(s(X_{1})=s(X_{n+1}))$
@@
$F_{Y}^{*}(s(X_{1}))\coloneqq\lim_{y\uparrow s(X_{1})}\,F_{Y}(y)+\theta_{1}\left(F_{Y}(s(X_{1}))-\lim_{y\uparrow s(X_{1})}\,F_{Y}(y)\right).$
@@
Fix $u\in(0,1)$ and suppose we want to test $\tilde{\mathcal{H}}_{0}:X_{1}\sim Q$ against $\tilde{\mathcal{H}}_{1}:X_{1}\sim R$, rejecting $\tilde{\mathcal{H}}_{0}$ whenever $F_{s(X_{n+1})}^{*}(s(X_{1}))\leq u$. Under $\tilde{\mathcal{H}}_{0}$, the cdf of $s(X_{1})$ is exactly $F_{s(X_{n+1})}$, so the randomized probability integral transform (Lemma F.1) gives that the Type I error of our test is $u$:
@@
The power of our test is $\mathbb{P}_{\tilde{\mathcal{H}_{1}}}\left(F_{s(X_{n+1})}^{*}(s(X_{1}))<u\right)$. By the generalized Neyman-Pearson lemma (see, for example, Theorem 3.2.1 and Section 5.9 of [lehmann2005testing]), the power is maximized by the choice $s=s^{*}$. Integrating the bound from 0 to 1 gives the desired result.
The usual Neyman-Pearson lemma [lehmann2005testing, Theorem 3.2.1 (ii)] states that any test $\phi$ of exact size $\alpha$ of the form $\phi(z)=\mathbf{1}_{s^{*}(z)\geq t_{\alpha}}$ is most powerful for testing $\mathcal{H}_{0}$ against $\mathcal{H}_{1}$ at level $\alpha$. Hence, we deduce that the choice $s^{*}$ maximizes the power of any level $\alpha$ test of $\mathcal{H}_{0}$ against $\mathcal{H}_{1}$. Furthermore, we know that $\phi_{\alpha}[s]$ is an exact test at size $\alpha$ ($\mathbb{E}[\phi_{\alpha}[s]]=\alpha$ when $Q=R$) which is independent of the choice of score $s$ [angelopoulos2024theoretical, Lemma 9.3]. Putting the pieces together, the last statement of the theorem follows. ∎
Appendix G The log-optimal e-variable uses the likelihood ratio score
While the reasoning in Theorem 5.1 yields that conformal p-values are optimized by the likelihood ratio score, it turns out to also be true that conformal e-values are also optimized by the same score. Indeed, in order to test exchangeability of $X_{1},\dots,X_{n+1}$ against the alternative that their joint distribution is given by $R^{n}\times Q$, one can show [ramdas2024hypothesis, Section 6.7.7] that the log-optimal e-variable is given by
@@ Appendix G The log-optimal e-variable uses the likelihood ratio score
Appendix H Convolutional neural network architecture
@@
Appendix I Additional methods for choosing the score function